OpenStudy (anonymous):

74) If the total number of diagonals of an n-gon is 77, then n = ________ ? a) 7 b) 8 c) 12 d) 14

OpenStudy (anonymous):

help me please. thanks

OpenStudy (anonymous):

Total no. of diagonals = nC2 - n. => nC2 - n = 77 (n(n-1) - 2n) = 154 n = 14

OpenStudy (anonymous):

It is nC2 - n because total diagonals can be understood as : Select any two points from all of them. Subtract n as A polygon has n sides and you dont want those in diagonals.